# DSOcc: Leveraging Depth Awareness and Semantic Aid to Boost Camera-Based 3D Semantic Occupancy Prediction

### v1

@@ Abstract
Camera-based 3D semantic occupancy prediction offers an efficient and cost-effective solution for perceiving surrounding scenes in autonomous driving. However, existing works rely on explicit occupancy state inference, leading to numerous incorrect feature assignments, and insufficient samples restrict the learning of occupancy class inference. To address these challenges, we propose leveraging Depth awareness and Semantic aid to boost camera-based 3D semantic Occupancy prediction (DSOcc). We jointly perform occupancy state and occupancy class inference, where soft occupancy confidence is calculated through non-learning method and multiplied with image features to make the voxel representation aware of depth, enabling adaptive implicit occupancy state inference. Rather than focusing on improving feature learning, we directly utilize well-trained image semantic segmentation and fuse multiple frames with their occupancy probabilities to aid occupancy class inference, thereby enhancing robustness. Experimental results demonstrate that DSOcc achieves state-of-the-art performance on the SemanticKITTI dataset among camera-based methods.
@@ 1 Introduction
The perception of surrounding scenes serves as the foundational part of autonomous driving. Compared to the LiDAR-based scheme [44, 65], camera-based perception [64, 8, 57] has attracted significant attention due to the lower cost of devices. To enable 3D perception for downstream tasks, camera-based 3D semantic occupancy prediction [32, 23, 62] represents the surrounding scene using a predefined voxel grid with semantic annotation. This task marks a significant advancement towards cheap, clear, efficient scene understanding for autonomous driving.
@@
For the occupancy state, two-stage methods [32, 42] and depth-query methods [23, 53] estimate a depth threshold to classify voxels as either occupied or unoccupied based on depth and voxel position. The issue arises when depth prediction or threshold estimation is inaccurate: since only occupied voxels are assigned features for subsequent classification, a truly occupied voxel may be misclassified as unoccupied, while its neighboring unoccupied voxel may incorrectly receive features. This phenomenon is not uncommon. Our analysis shows that in [32], nearly 40% of occupied voxels are misclassified as unoccupied. To mitigate accumulated errors, it is crucial to reduce the heavy reliance on depth prediction or threshold estimation. For the occupancy class, the prediction accuracy relies on the mapping effect from image features to occupancy classes. To this end, previous methods have incorporated various modules to learn geometric features [56], instance cues [23], and semantic guidance [42]. However, a critical issue overlooked is that they are trained on limited datasets caused by high annotation costs. For instance, SemanticKITTI [3] and SSCBenchKITTI-360 [33] contain only 3,834 and 8,487 training samples, respectively. Within this fact, effectively utilizing image features is akin to ’dancing in shackles’, and the robustness of occupancy class prediction are constrained in this shackle.
@@ 2 Related Works
For camera-based 3D semantic occupancy prediction, dimensional transformation is a crucial component. MonoScene [6] firstly introduced a 2D-to-3D feature projection. OccFormer [62] leverages local and global transformers to perform projection along the horizontal plane. SurroundOcc [54] employed spatial projection to lift multi-scale features. COTR [41] introduced an explicit-implicit transformation for compact 3D representation. ViewFormer [28] proposed a learning-first view attention mechanism to aggregate multi-view features. Following dimensional transformation, various scenes representations have been proposed, beyond voxels. TPVFormer [20] introduced three perpendicular planes to learn image features. GaussianFormer [22] used sparse 3D semantic Gaussians to represent dynamic scenes. OSP [48] proposed a point-based representation by treating scenes as point clouds.
@@
2D Image Representation. Besides image features, camera-based 3D semantic occupancy prediction also incorporates information from other representations, such as depth, semantic segmentation, and instance segmentation. Depth estimation [13, 10] is a fundamental task in camera-based autonomous driving and can be categorized into relative depth [58, 16] and absolute depth [21, 17, 45]. Using monocular [30, 55, 12] or stereo [4, 49] image inputs, absolute depth estimation provides the real distance between objects and the camera. Absolute depth information has been applied in camera-based 3D semantic occupancy prediction to resolve projection ambiguity in the frontal direction [32, 53], thereby helping to determine occupancy state.
Semantic segmentation [38] predicts a semantic label for each image pixel, while instance segmentation [31] distinguishes objects within a class based on it. Cityscapes [9], KITTI [1], BDD100K [60], and nuScenes [5] provide diverse driving scenes under varying conditions. By training on these datasets, many classic models, such as FCNs [38], DeepLab [7], and SegNet [2], achieve robust and efficient segmentation for real-time autonomous driving. Recently, universal segmentation models [25, 29, 51, 35] aim to improve scalability by training on mixed datasets.
@@ 3.1 Overview
DSOcc aims to enhance the inference capacity of the occupancy state (whether) and the occupancy class (which) in 3D semantic occupancy prediction. It utilizes a pre-trained image encoder to extract image features , and employs a well-trained depth predictor and semantic segmentation model to infer depth maps and semantic segmentation map . To address the challenges of ambiguous spatial correspondences between voxels and images in occupancy state inference, we lift 2D image features into depth-aware image feature voxels for each frame and obtain by computing a soft occupancy confidence and mapping image features into a dense voxel through backward projection. To aid the occupancy class inference with semantic segmentation map, we lift semantic segmentation maps to a semantic-aided voxel by integrating multi-frame one-hot representations with depth-aware confidence. Concurrently, we propose a novel voxel fusion module that fuses and to obtain by utilizing global interaction through deformable attention.
We will discuss the depth-aware image feature voxel and the semantic-aided voxel in Sec. 3.2 and Sec. 3.3, respectively, and introduce the voxel fusion module in Sec. 3.4.
Depth Predictor. It estimates the absolute depth between objects and ego vehicle, conditioned on camera images and calibrations inputs. Specifically, we leverage the well-trained MobileStereoNet [46], similar to previous approaches [23, 32]. This model is lightweight and well-suited for real-time autonomous driving applications.
Semantic Segmentation Model. It demonstrates strong robustness and high prediction accuracy when trained on large datasets. We leverage this capability to improve the occupancy classe inference. Specifically, we employ MSeg [25] and adapt its universal class classification to the specific class presets required for this task.
@@ 3.3 Semantic-Aided Voxel
The core challenge in 3D semantic occupancy prediction is to identify the class of each voxel based on image features. However, due to limitations in model capacity and the domain represented by datasets, prediction models face challenges such as class imbalance and poor class discrimination, as illustrated in Fig. 4c. In contrast, image semantic segmentation benefit from larger and more diverse datasets, enabling segmentation models to exhibit better robustness. Fundamentally, both image semantic segmentation and 3D semantic occupancy prediction share the same objective, and the former class distribution closely resembles the latter ground truth, as illustrated in Fig. 4a and Fig. 4b. Therefore, we aim to use the former capacity to the latter task to aid occupancy class inference.
We establish a semantic-aided voxel to collect segmentation information from using the projection relationship defined as Equ. 1. As illustrated in Fig. 4d, the same voxel may collect different classes from different frames due to the projection shift between adjacent frames. Since the discrete segmentation is a high-level feature, effectively utilizing segmentation information from multiple frames is a key challenge.
@@ 4.1 Dataset
Dataset. We evaluate DSOcc on the SemanticKITTI [3] and SSCBenchKITTI-360 [33] datasets, which provide 3D semantic occupancy annotations from the LiDAR sweeps of the KITTI Odometry Benchmark [14]. Both datasets focus on a scene of , with a voxel size of , resulting in a voxel grid. SemanticKITTI includes 22 outdoor driving scenarios, with 10 sequences for training, 1 for validation, and 11 for testing. It annotates voxels into 20 semantic classes (19 semantics + 1 free) and has resolution camera images. SSCBenchKITTI-360 consists of 9 outdoor driving scenarios, with 9 sequences for training, 1 for validation, and 1 for testing. It has 19 class labels (18 semantics + 1 free) and resolution camera images.
Metric. Following SSC, we evaluate model performance in terms of intersection over union (IoU) and mean intersection over union (mIoU) across 19 classes. IoU measures overall scene completion quality, while mIoU assesses the quality of semantic segmentation for each class. We use mIoU as the primary metric.
@@ 4.2 Implementation Details
We crop images into size , and use ResNet50 [18] to extract its feature map. All experiments were conducted on 2 NVIDIA A6000 GPUs with a batch size of 1. The sizes of depth-aware image feature voxel and semantic-aided voxel are set to as the final prediction. The DA used in the voxel fusion module has 1 layer and 8 sampling points. We optimize DSOcc using the AdamW optimizer [39] with an initial learning rate of and a weight decay of for a total of 40 epochs. Following the method in [32, 52, 42, 53], we propose two versions of the DSOcc model: DSOcc-T, which takes the current and previous four frames as inputs, and DSOcc-S, which takes only the current frame as input.
@@ 4.3 Main Results
Qualitative Comparisons. We present the qualitative comparison results on the SemanticKITTI validation set, the SemanticKITTI hidden test set, and the SSCBench-KITTI-360 test set in Table 1, Table 2, and Table 3, respectively. On the SemanticKITTI dataset, DSOcc-T achieves the highest mIoU of 18.02 (+2.14) and 16.90 (+0.17) on the validation and hidden test sets, respectively, outperforming 12 and 9 classes out of the total 19 classes. DSOcc-S also demonstrates state-of-the-art performance among S-version models, with mIoU scores of 14.59 (+0.04) and 14.96 (+0.95) on the validation and hidden test sets, respectively. On the SSCBench-KITTI-360 test set, DSOcc maintains competitive performance, achieving 17.51 mIoU for S-version, demonstrating the robustness of our method.
@@
Qualitative Comparisons. We present the qualitative comparison results on the SemanticKITTI validation set in Fig. 6. DSOcc can perceive small objects far from the ego vehicle, such as bicyclists and poles, and can clearly distinguish every vehicle in a convoy. In complex scenes, such as corners and shops, DSOcc demonstrates excellent discrimination ability between different classes within a local region. This highlights its long-range and class-balanced scene perception, enhancing the reliability and robustness of the entire autonomous driving system.
@@ 4.4 Ablation Studies
All ablation studies were conducted on the SemanticKITTI val set for DSOcc-T. We focus on four aspects: scene range, input representation, feature processing, and fusion mechanism.
@@
Input Representation. To evaluate the impact of different input representations, we assess four input types: image features in 2D form (Img. 2D), image features in 3D voxel form (Img. 3D), depth, and semantic segmentation map (Sem.). We use the projection-learning module from [32] for the Img. 2D case. As shown in Table 5, aggregating image features into a dense voxel grid facilitates the learning of spatial geometric relationships, resulting in a 1.22 mIoU improvement. Depth can enhance scene completion, yielding a 0.22 mIoU increase. We report that depth-aware image feature voxels and semantic-aided voxels achieve 15.00 and 11.16 mIoU, respectively, with standard deviations of 15.09 and 12.76 across all classes. The semantic-aided voxel demonstrates more accurate predictions, especially for low-frequency classes. For example, the truck class (0.16%) achieves an IoU of 10.28 vs 3.99. It demonstrates that image semantic segmentation contributes to class-balanced predictions, improving mIoU to 18.02 for the fusion case.
@@
Feature Processing. To evaluate different depth embedding ways for establishing depth-aware image feature voxel, we compare baseline with two alternatives: depth-concatenation and depth-truncation. They concatenate depth along the channel dimension or explicitly set a depth threshold to determine occupancy status. As depicted in Table 6, these methods either increase the burden on feature fusion learning or restrict feature collection, decreasing mIoU by 5.65 and 5.63, respectively. To assess the impact of multi-frame processing in establishing semantic-aided voxel, we compare the baseline with two strategies: using only a single frame (w/t multi. sem.) and concatenating multiple frames (w/t sem. concat.). These approaches either miss potential class information or fail to identify the importance of individual frames.
@@
Fusion Mechanism. To evaluate the design of the voxel fusion module, we compare it with five alternatives: using Conv3D layers (Conv.), using Swin Transformer (Swin.), using as the query with DA (Single DA =img.), using as the query with DA (Single DA. =sem.), and using dual DA with 2 layers (DA 2 layers). As depicted in Table 7, Conv. and Swin. encounter challenges in capturing long-range dependencies, decreasing mIoU by 1.27 and 1.49, respectively. Removing either interaction in baseline also results in a drop in mIoU. The 2-layer version does not yield any prediction increase but increases inference time to 283.76. In a nutshell, our voxel fusion module strikes a balance between precision and efficiency.
| Inf. Time (ms) | IoU | mIoU | |
| Conv. | 0.16 | 40.55 | 16.75 (-1.27) |
| Swin. | 0.67 | 41.30 | 16.53 (-1.49) |
| Single DA (=Img.) | 134.52 | 44.69 | 16.94 (-1.08) |
| Single DA (=Sem.) | 134.12 | 42.91 | 17.88 (-0.14) |
| 2 Layers DA | 283.76 | 45.16 | 17.48 (-0.54) |
| Ours | 177.07 | 44.41 | 18.02 |
@@ 5 Conclusion
In this paper, we propose DSOcc, which leverages depth-aware image feature voxel and semantic-aided voxel to boost 3D semantic occupancy prediction. DSOcc embeds soft occupancy confidence into image features to enable adaptive implicit occupancy state inference, and utilizes image semantic segmentation information to aid occupancy class inference, thereby improving robustness. Experimental results demonstrate that DSOcc achieves state-of-the-art performance on the SemanticKITTI dataset among camera-based methods.
@@ Appendix A Occupancy Status Inference for Two-stage Method
We evaluate the Stage-1 module in Voxformer [32] on the SemanticKITTI validation dataset [3]. The occupancy inference results are depicted in Table 8. While the result of appear promising, it is important to note that occupied voxels represent a small proportion of all voxels ( = 14.60%). As a result, 38.14% of occupied voxels are incorrectly inferred as unoccupied. These false negatives do not collect any image features in the subsequent stage, leading to numerous inference errors.
@@ Appendix B The Class Mapping in Image Semantic Segmentation
We use MSeg [25] to predict semantic segmentation maps from camera images, which provides a universal class classification. To align these predictions with the class settings of 3D semantic occupancy prediction, we perform class mapping during the preprocessing step. The class mappings for SemanticKITTI and SSCBench-KITTI-360 [33] are shown in Tables 9 and 3, respectively. This demonstrates that strict alignment in class mapping is not required. Even when some classes are missing in the SemanticKITTI dataset, the semantic-aided voxel still provides sufficient semantic guidance for 3D semantic occupancy prediction.

### v3

@@ Abstract
Camera-based 3D semantic occupancy prediction offers an efficient and cost-effective solution for perceiving surrounding scenes in autonomous driving. However, existing works rely on explicit occupancy state inference, leading to numerous incorrect feature assignments, and insufficient samples restrict the learning of occupancy class inference. To address these challenges, we propose leveraging Depth awareness and Semantic aid to boost camera-based 3D semantic Occupancy prediction (DSOcc). We jointly perform occupancy state and occupancy class inference, where soft occupancy confidence is calculated by non-learning method and multiplied with image features to make voxels aware of depth, enabling adaptive implicit occupancy state inference. Instead of enhancing feature learning, we directly utilize well-trained image semantic segmentation and fuse multiple frames with their occupancy probabilities to aid occupancy class inference, thereby enhancing robustness. Experimental results demonstrate that DSOcc achieves state-of-the-art performance on the SemanticKITTI dataset among camera-based methods.
@@ 1 Introduction
The perception of surrounding scenes serves as the foundational part of autonomous driving. Compared to the LiDAR-based scheme [41, 57], camera-based scheme [56, 7] has attracted significant attention due to the lower cost of devices. To achieve 3D perception, camera-based 3D semantic occupancy prediction [27, 54] represents the surrounding environment using a predefined voxel grid, where each voxel is annotated with a semantic label. This task marks a significant advancement towards cheap, clear, efficient scene understanding for autonomous driving.
@@
For the occupancy state, two-stage methods [27, 38] and depth-query methods [19, 47] estimate a depth threshold to classify voxels as either occupied or unoccupied based on depth and voxel position. The issue arises when depth prediction or threshold estimation is inaccurate: since only occupied voxels are assigned features for subsequent classification, a truly occupied voxel may be misclassified as unoccupied, while its neighboring unoccupied voxel may incorrectly receive features. This phenomenon is common. Our analysis shows that in [27], nearly 40% of occupied voxels are misclassified as unoccupied (see Appendix A). To mitigate accumulated errors, it is crucial to reduce the heavy reliance on depth prediction or threshold estimation. For the occupancy class, the prediction accuracy relies on the mapping effect from image features to occupancy classes. To this end, previous methods have incorporated various modules to learn geometric features [50], instance cues [19], and semantic guidance [38]. However, a critical issue overlooked is that they are trained on limited datasets. For instance, SemanticKITTI [2] and SSCBenchKITTI-360 [28] contain only 3,834 and 8,487 training samples, respectively. Within this fact, effectively utilizing image features is akin to ’dancing in shackles’, and the occupancy class prediction are constrained in this shackle.
@@ 2 Related Works
For camera-based 3D semantic occupancy prediction, dimensional transformation is a crucial component. MonoScene [5] firstly introduced a 2D-to-3D feature projection. OccFormer [54] leverages local and global transformers to perform projection along the horizontal plane. SurroundOcc [49] employed spatial projection to lift multi-scale features. COTR [37] introduced an explicit-implicit transformation for compact 3D representation. ViewFormer [24] proposed a learning-first view attention mechanism to aggregate multi-view features.
@@
2D Image Representation. Besides image features, camera-based 3D semantic occupancy prediction also incorporates information from other inputs, such as depth, semantic segmentation, and instance segmentation. Depth estimation is a fundamental task in camera-based autonomous driving and can be categorized into relative depth [51, 13] and absolute depth [17, 10]. With monocular [25, 11] or stereo [3] images, absolute depth estimation provides the real distance between objects and the camera. Absolute depth information has been applied in camera-based 3D semantic occupancy prediction to resolve projection ambiguity in the frontal direction [27, 47], thereby helping to determine occupancy state.
Semantic segmentation [34] predicts a semantic label for each image pixel, while instance segmentation [26] distinguishes objects within a class based on it. Cityscapes [8], KITTI [1], and nuScenes [4] provide diverse driving scenes and conditions. By training on these datasets, many classic models, such as FCNs [34], and DeepLab [6], achieve robust and efficient segmentation for real-time application. Recently, universal segmentation models [21, 31] aim to improve scalability by training on mixed datasets.
@@ 3.1 Overview
DSOcc aims to enhance the inference capacity of the occupancy state (whether) and the occupancy class (which) in 3D semantic occupancy prediction. It utilizes a pre-trained image encoder to extract image features , and employs a well-trained depth predictor and semantic segmentation model to infer depth maps and semantic segmentation map . To address the challenges of ambiguous spatial correspondences between voxels and images in occupancy state inference, we lift 2D image features into depth-aware image feature voxels for each frame and obtain by computing a soft occupancy confidence and backward project image features into a dense voxel. To aid the occupancy class inference, we lift semantic segmentation maps to a semantic-aided voxel by integrating multi-frame one-hot representations with occupancy probabilities. Concurrently, we propose a novel voxel fusion module that fuses and to obtain by utilizing global interaction through deformable attention.
Depth Predictor. It estimates the absolute depth between objects and ego vehicle, conditioned on camera images and calibrations. Specifically, we leverage the well-trained MobileStereoNet [42], similar to previous method [19, 27]. This model is lightweight and well-suited for real-time applications.
Semantic Segmentation Model. It is pretrained on large-scale datasets, and exact class alignment with our task is not required. Specifically, we employ MSeg [21] and adapt its universal class classification to the specific class presets required for this task (see Appendix B).
@@ 3.3 Semantic-Aided Voxel
The core challenge in 3D semantic occupancy prediction is to identify the class of each voxel based on image features. However, due to limitations in model capacity and the dataset scale, prediction models face challenges such as class imbalance and poor class discrimination, as illustrated in Fig. 4c. In contrast, image semantic segmentation benefit from larger and more diverse datasets, achieving greater robustness. Fundamentally, both image semantic segmentation and 3D semantic occupancy prediction share the same objective, and the former class distribution closely resembles the latter ground truth, as illustrated in Fig. 4a and Fig. 4b. Therefore, we aim to use the former capacity in the latter task to aid occupancy class inference.
We establish a semantic-aided voxel to collect segmentation information from using the projection relationship defined as Equ. 1. As illustrated in Fig. 4d, the same voxel may collect different classes from different frames due to the projection shift between adjacent frames. Since the results are discrete, effectively utilizing segmentation information from multiple frames is a key challenge.
@@ 4.1 Dataset
Dataset. We evaluate DSOcc on the SemanticKITTI [2] and SSCBenchKITTI-360 [28] datasets. Both datasets focus on a scene of , with a voxel size of . SemanticKITTI includes 20 class labels (19 semantics + 1 free) , whereas SSCBenchKITTI-360 includes 19 class labels (18 semantics + 1 free).
Metric.Following SSC, we measure performance using intersection over union (IoU) and mean IoU (mIoU) over 19 classes. IoU reflects overall scene completion quality, while mIoU the quality of semantic segmentation for each class. We report mIoU as our primary metric.
@@ 4.2 Implementation Details
We crop images into size , and use ResNet50 [14] to extract its feature map. All experiments were conducted on 2 NVIDIA A6000 GPUs with a batch size of 1. The DA used in the voxel fusion module has 1 layer and 8 sampling points. We optimize DSOcc using the AdamW optimizer [35] with an initial learning rate of and a weight decay of for a total of 40 epochs. Following the method in [27, 46, 38, 47], we propose two versions of the DSOcc model: DSOcc-T, which takes the current and previous four frames as inputs, and DSOcc-S, which takes only the current frame as input.
@@ 4.3 Main Results
Qualitative Comparisons. We present the qualitative comparison results on the SemanticKITTI validation set, the SemanticKITTI hidden test set, and the SSCBench-KITTI-360 test set in Table 1, Table 2, and Table 3, respectively. For a more comprehensive evaluation, additional comparisons on Occ3D-NuScenes [45] are included in Appendix C.3. On the SemanticKITTI dataset, DSOcc-T achieves the highest mIoU of 18.02 (+2.14) and 16.90 (+0.17) on the validation and hidden test sets, respectively, outperforming 12 and 9 classes out of the total 19 classes. DSOcc-S also demonstrates state-of-the-art performance among S-version models, with mIoU scores of 14.59 (+0.04) and 14.96 (+0.95) on the validation and hidden test sets, respectively. On the SSCBench-KITTI-360 test set, DSOcc maintains competitive performance, achieving 17.51 mIoU for S-version, demonstrating the robustness of our method.
@@
Qualitative Comparisons. We present the qualitative comparison results on the SemanticKITTI validation set in Fig. 6, with additional examples provided in Appendix C.2. DSOcc can perceive small objects far from the ego vehicle, such as bicyclists and poles, and can clearly distinguish every vehicle in a convoy. In complex scenes, such as corners and shops, DSOcc demonstrates excellent discrimination ability within a local region. This highlights its long-range and class-balanced scene perception, enhancing the reliability and robustness of autonomous driving.
@@ 4.4 Ablation Studies
All ablation studies were conducted on the SemanticKITTI val set for DSOcc-T. We focus on four aspects: scene range, input representation, and feature processing. Additional ablation studies are provided in Appendix C.1.
@@
Input Representation. To evaluate the impact of different input representations, we assess four input types: image features in 2D form (Img. 2D), image features in 3D voxel form (Img. 3D), depth, and semantic segmentation map (Sem.). We use the projection-learning module from [27] for the Img. 2D case. As shown in Table 5, Img. 3D improves spatial geometry learning by 0.17 mIoU. Incorporating depth enhances scene completion, yielding a 2.90 IoU gain. We report that depth-aware image feature voxels and semantic-aided voxels achieve 15.00 and 11.16 mIoU, respectively, with standard deviations of 15.09 and 12.76 across all classes. The semantic-aided voxel achieves higher precision on low-proportion classes. For example, the truck class (0.16 %) improves from 3.99 to 10.28. It shows that semantic-aided voxels promote class-balanced predictions.
@@
Feature Processing. To assess depth embedding methods, we compare the baseline against two variants: depth-concatenation and depth-truncation. They appends the depth channel or explicitly set a depth threshold to determine occupancy status. As depicted in Table 6, these methods either increase the burden on feature fusion learning or restrict feature collection, decreasing mIoU by 5.65 and 5.63, respectively. To assess the impact of multi-frame processing in establishing semantic-aided voxel, we compare the baseline with two strategies: using only a single frame (w/t multi. sem.) and concatenating multiple frames (w/t sem. concat.). These approaches either miss potential class information or fail to identify the importance of individual frames.
@@ 5 Conclusion
In this paper, we propose DSOcc, which leverages depth-aware image feature voxel and semantic-aided voxel to boost 3D semantic occupancy prediction. DSOcc embeds soft occupancy confidence into image features to enable adaptive implicit occupancy state inference, and utilizes image semantic segmentation to aid occupancy class inference, thereby improving robustness. Experimental results demonstrate that DSOcc achieves state-of-the-art performance on the SemanticKITTI dataset among camera-based methods.
@@ Appendix A Occupancy Status Inference for Two-stage Method
We evaluate the Stage-1 module in Voxformer [27] on the SemanticKITTI validation dataset [2]. The occupancy inference results are depicted in Table A1. While the result of appear promising, it is important to note that occupied voxels represent a small proportion of all voxels ( = 14.60%). As a result, 38.14% of occupied voxels are incorrectly inferred as unoccupied. These false negatives do not collect any image features in the subsequent stage, leading to numerous inference errors.
@@ Appendix B The Class Mapping in Image Semantic Segmentation
We use MSeg [21] to predict semantic segmentation maps from camera images, which provides a universal class classification. To align these predictions with the class settings of 3D semantic occupancy prediction, we perform class mapping during the preprocessing step. The class mappings for SemanticKITTI and SSCBench-KITTI-360 [28] are shown in Tables B1 and B2, respectively. This demonstrates that strict alignment in class mapping is not required. Even when some classes are missing in the SemanticKITTI dataset, the semantic-aided voxel still provides sufficient semantic guidance for 3D semantic occupancy prediction.